OpenStudy (anonymous):

Write the logarithm as an exponent equation: log12 4=x, please show how you did it so i could see if i did it right? ^^;

OpenStudy (valpey):

You are asking:?\[\large \log_{12}{4}=x\]\[\large 12^{\log_{12}{4}}=12^x\]\[\large 4=12^x\]

OpenStudy (lgbasallote):

just remember this rule \[\huge \log_a b = c \implies a^c = b\]